The vectors from the origin to the points A, B,


C and D are A= i + j + k


B = 2i +3j


C= 3i + 5j -2k


D = k- j


Show that the lines AB and CD are parallel and find the ratios of their lengths

1
Expert's answer
2022-01-13T09:20:49-0500

Find how the lines can be described in unit vectors:


"AB=\vec B-\vec A=(2\hat i+3\hat j)-(\hat i+\hat j+\hat k)=\\=\hat i+2\hat j-\hat k,\\\space\\
CD=\vec D-\vec C=(-\hat j+\hat k)-(3\hat i+5\hat j-2\hat k)=\\=-3\hat i-6\hat j+3\hat k."


Find the dot product:


"AB·CD=(\hat i+2\hat j-\hat k)·(-3\hat i-6\hat j+3\hat k)=\\
=-3-12-3=-18."


Find the length of individual line:


"|AB|=\sqrt{1^2+2^2+(-1)^2}=\sqrt 6.\\
|CD|=\sqrt{(-3)^2+(-6)^2+3^2}=\sqrt54.\\\space\\
\dfrac{|CD|}{|AB|}=3."

Since the dot product is


"AB·CD=|AB|·|CD|\cos\theta,"

we can find the angle "\theta" between the lines to establish the configuration of the lines relative to each other:


"\theta=\arccos\dfrac{AB·CD}{|AB|·|CD|}=180°."


As we see, the lines are parallel to each other.